-2/5m= -6/7
solve and check

Answer by edjones(8007)

-2m/5= -6/7 I assume this is the problem.
2m/5=6/7
35*2m/5=35*6/7
14m=30 No fractions.
14m/14=30/14
m=15/7
.
Check:
-2*15/7 /5 = -6/7
-30/7 * 1/5 = -6/7 True
.
Ed